# A machine-learning photometric classifier for massive stars in nearby galaxies

### v1

@@ Abstract
Methods. We combined Spitzer photometry and Pan-STARRS1 optical data to classify point sources in 26 galaxies within 5 Mpc, and a metallicity range 0.07-1.36 Z. Gaia DR3 astrometry was used to remove foreground sources. Classifications are derived using a machine-learning model developed by Maravelias et al. (2022).
Results. We report classifications for 1,147,650 sources, with 276,657 sources ($\sim 24\%$) being robust. Among these are 120,479 Red Supergiants (RSGs; $\sim 11\%$). The classifier performs well even at low metallicities ($\sim 0.1$ Z) and distances under 1.5 Mpc, with a slight decrease in accuracy beyond $\sim 3$ Mpc due to Spitzer ’s resolution limits. We also identified 21 luminous RSGs ($\textrm{log}(L/L_{\odot})\,$$\geq 5.5$), 159 dusty Yellow Hypergiants in M31 and M33, as well as 6 extreme RSGs ($\textrm{log}(L/L_{\odot})\,$$\geq 6$) in M31, challenging observed luminosity limits. Class trends with metallicity align with expectations, though biases exist.
Conclusions. This catalog serves as a valuable resource for individual-object studies and James Webb Space Telescope target selection. It enables follow-up on luminous RSGs and Yellow Hypergiants to refine our understanding of their evolutionary pathways. Additionally, we provide the largest spectroscopically confirmed catalog of massive stars and candidates to date, comprising 5,273 sources (including $\sim 330$ other objects).
@@
Stars: massive – Stars: mass-loss – Stars: evolution – Galaxies: individual: WLM, NGC 55, IC 10, M31, NGC 247, NGC 253, NGC 300, IC 1613, M33, Phoenix Dwarf, NGC 1313, NGC 2366, NGC 2403, M81, Sextans B, NGC 3109, NGC 3077, Sextans A, NGC 4214, NGC 4736, NGC 4826, M83, NGC 5253, NGC 6822, Pegasus DIG, NGC 7793 – Methods: statistical – Catalogs
@@ 1 Introduction
One of the main goals of James Webb Space Telescope (JWST) is to study galaxies over cosmic time, from the early Universe to now. Their light though is a combination of various components, one of which is their stellar content. And, although rare in absolute numbers, massive stars have an important contribution. Through their feedback, whether from strong stellar winds or explosive supernovae, massive stars play a critical role in enriching and shaping the environments of their host galaxies. This is especially important in the early Universe, when metallicity was extremely low and the formation and evolution of such stars is still not well understood. The only way to gain insight into these objects is by examining these populations in nearby low metallicity galaxies. Resolved population studies in such galaxies are possible in the Local Group (e.g. Sextans A; Lorenzo et al. 2022, LMC; Vink et al. 2023, SMC; Shenar et al. 2024) but challenging at larger distances. Therefore, we lack well-explored populations of massive stars at these metalliclities.
@@
The main goal of the ASSESS [^1] http://assess.astro.noa.gr/ (”Episodic Mass Loss in Evolved Massive stars: Key to Understanding the Explosive Early Universe”) project (Bonanos et al., 2024) was to investigate the role of episodic mass loss in the evolution of massive stars (see e.g. Yang et al. 2023; Antoniadis et al. 2024, 2025; de Wit et al. 2024; Munoz-Sanchez et al. 2024a, b; Zapartas et al. 2024). A large number of sources with secure classifications was needed to explore its importance across various metallicity environments, therefore we set up both an observing campaign to acquire spectra for a large number of sources (see e.g. de Wit et al., 2023; Bonanos et al., 2024; de Wit et al., subm.) as well as a machine-learning approach (Maravelias et al. 2022; henceforth Paper I). We developed a classifier that uses optical and IR photometry to select dusty, mass losing, evolved massive stars and classify Spitzer detected point sources into different broad classes. The purpose of this classifier was to predict the classes for approximately 1.2 M sources from 26 galaxies within 5 Mpc and spanning a metallicity range (0.07-1.36 Z), creating the largest point-source catalog with spectral-type classifications for and beyond the Local Group. In Paper I, we provide the method and explore the classifier’s prediction accuracy [^2] Code and other material availabe at https://github.com/gmaravel/pc4mas., while the current paper presents the results of the classifier’s application and the corresponding catalog of sources.
@@
In Section 2 we present the data collection and processing, the selection process of the best candidates (based on the results from the machine-learning classifier), as well as the collection of all sources with known spectral classification from literature. In Section 3 we describe the catalog we compiled along with statistics regarding the number of objects per class, and present color-magnitude diagrams (CMDs). In Section 4 we provide a comparison of the spectral types we predicted with the ones derived from the literature and discuss the performance of the classifier. We present the populations as a function of metallicity and explore luminous Red Supergiants (RSGs) and dusty Yellow Supergiants (YSGs) found in our sample. Finally, in Section 5 we summarize and conclude our work.
@@ 2 Data collection and processing
In the following sections, we describe our sample and the steps we followed to build our catalogs as well as to remove foreground sources. Although this has been extensively discussed in Paper I, we provide here a short overview, along with adjustments performed mainly regarding the foreground cleaning due to the release of Gaia DR3.
@@ 2.1 Galaxy sample
In Paper I we used M31 and M33 for training the classifier, and WLM, Sextans A, and IC 1613 for testing it. In this work, we present the results of applying the classifier to the whole sample of galaxies included in the ASSESS project (Bonanos et al., 2024) (except for the LMC and SMC which were treated separately, e.g. Yang et al. 2019, 2020, 2021, 2023; de Wit et al. 2023; Antoniadis et al. 2024; Munoz-Sanchez et al. 2024a). This is presented in Table 1 along with the basic properties of the galaxies, including their names, coordinates, types, radii (which correspond to the sizes of the galaxy based on visual inspection that was used to match the catalogs), distances, and metallicity.
Table 1: Galaxies examined in this work along with some basic properties.
@@
(hh:mm:ss) (dd:mm:ss) (′) (Mpc) ($Z_{\sun}$)
WLM 00:01:58 --15:27:39 SB(s)m: sp 9 0.95$\pm$0.01 0.131
NGC 55 00:14:54 --39:11:48 SB(s)m? edge-on 21 1.98$\pm$0.02 0.312
IC 10 00:20:17 +59:18:14 dIrr IV/BCD 6 0.78$\pm$0.04 0.473
M31 00:42:44 +41:16:09 SA(s)b LINER 105 0.75$\pm$0.02 1.364
NGC 247 00:47:09 --20:45:37 SAB(s)d 14 3.56$\pm$0.03 0.545
NGC 253 00:47:33 --25:17:18 SAB(s)c 21 3.61$\pm$0.03 0.836
NGC 300 00:54:53 --37:41:04 SA(s)d 15 1.94$\pm$0.04 0.417
IC 1613 01:04:48 +02:07:04 IB(s)m 14 0.72$\pm$0.01 0.168
M33 01:33:51 +30:39:37 SA(s)cd HII 30 0.85$\pm$0.02 0.659
Phoenix Dwarf 01:51:06 --44:26:41 IAm 8 0.43$\pm$0.01 0.0710
NGC 1313 03:18:16 --66:29:54 SB(s)d 8 4.21$\pm$0.06 0.3511
NGC 2366 07:28:55 +69:12:57 IB(s)m 6 3.21$\pm$0.04 0.1612
NGC 2403 07:36:51 +65:36:09 SAB(s)cd 14 3.13$\pm$0.06 0.5613
M81 09:55:33 +69:03:55 SA(s)ab 18 3.61$\pm$0.22 $0.60^{14}$
Sextans B 10:00:00 +05:19:56 IB(s)m 5 1.38$\pm$0.08 0.1015
NGC 3109 10:03:07 --26:09:35 SB(s)m edge-on 13 1.37$\pm$0.08 0.1216
NGC 3077 10:03:19 +68:44:02 I0 pec 5 3.75$\pm$0.11 0.8917
Sextans A 10:11:01 --04:41:34 IBm 4 1.41$\pm$0.05 0.0715
NGC 4214 12:15:39 +36:19:37 IAB(s)m 7 2.82$\pm$0.09 0.3218
NGC 4736 12:50:53 +41:07:14 (R)SA(r)ab 11 4.34$\pm$0.08 0.4819
NGC 4826 12:56:44 +21:40:59 (R)SA(rs)ab 7 4.36$\pm$0.06 0.2220
M83 13:37:01 --29:51:56 SAB(s)c 10 4.79$\pm$0.09 0.9721
NGC 5253 13:39:56 --31:38:24 Im pec 4 3.55$\pm$0.21 0.3722
NGC 6822 19:44:58 +14:48:12 IB(s)m 10 0.46$\pm$0.01 0.3023
Pegasus DIG 23:28:36 +14:44:35 dI 8 0.95$\pm$0.03 0.1024
NGC 7793 23:57:50 +32:35:28 SA(s)d 8 3.39$\pm$0.06 0.4825
@@
1Urbaneja et al. (2008); 2Hartoog et al. (2012); 3Cosens et al. (2024); 4Zurita & Bresolin (2012, range 1.05-1.66 $Z_{\sun}$); 5Davidge 2021, similar to LMC, 0.43 Z from Hunter et al. (2007) and M33 - see below; 6Spinoglio et al. (2022); 7Kudritzki et al. (2008); 8Bresolin et al. (2007); 9U et al. (2009, range 0.3-1.0 $Z_{\sun}$); 10Ross et al. (2015); 11Hadfield & Crowther (2007); 12Thuan & Izotov (2005); 13Bresolin et al. (2022, range 0.35-0.77 $Z_{\sun}$); 14Arellano-Córdova et al. (2016, range 0.37-0.78 $Z_{\sun}$); 15Kniazev et al. (2005, range 0.07-0.14 $Z_{\sun}$); 16Hosek et al. (2014); 17Storchi-Bergmann et al. (1994); 18Pilyugin et al. (2015); 19Moustakas & Kennicutt (2006); 20Kang et al. (2020, from iron abundance, range 0.03-0.40 $Z_{\sun}$); 21Hernandez et al. (2019, range 0.35-1.58 $Z_{\sun}$); 22Monreal-Ibero et al. (2012); 23Patrick et al. (2015); 24Skillman et al. (1997, range 0.06-0.20 $Z_{\sun}$) 25Della Bruna et al. (2021).
333 \$a\$\$a\$footnotetext: Derived from NASA/IPAC Extragalactic Database (NED). \$b\$\$b\$footnotetext: From Tully et al. (2023), except for M81, Sextans B, NGC 3109, and NGC 5253 which were derived from Tully et al. (2013). \$b\$\$b\$footnotetext: Metallicities are based on young clusters or massive stars or H ii regions, when available. An average value was reported in cases where multiple measurements or radial gradients exist. For solar abundance we used 12+log(O/H) = 8.69 (Asplund et al., 2009).
Figure 1: Example of the fitting processing and foreground removal in NGC 2403. Top panel: Gaia sources where the green ellipse defines the boundary selected for the galaxy. Middle panel: Fitting the parallax data for all sources within NGC 2403. The foreground contribution (dark gray dotted line) is scaled according to the densities of sources inside and outside the galaxy, while the purple dashed line shows the (Gaussian) contribution of galactic sources, and the blue line indicates the total fit. The vertical black dashed lines show the $3\sigma$ limits defined by the Gaussian properties of the galactic sources. Bottom panel: Similar to parallax, but for proper motion in declination (see Section 2.3 for more details).
@@ 2.2 Surveys used and data processing
Since IR photometry has been proven successful in identifying evolved massive stars (Britavskiy et al., 2014, 2015, 2019; Kourniotis et al., 2017) and those with dusty environments (Bonanos et al., 2009, 2010, 2024), we based our catalogs on pre-compiled point-source catalogs from the Spitzer Space Telescope (Boyer et al. 2015; Khan et al. 2015; Khan 2017; Williams & Bonanos 2016 - see Table 2 for details; the Spitzer column provides the number of sources within the radius as presented in Table 1). The mid-IR photometry ($\rm 3.6\,\mu m$, $4.5\,\mu\rm m$, $5.8\,\mu\rm m$, $8.0\,\mu\rm m$, $24\,\mu\rm m$ [^4] We are using the [3.6], [4.5], [5.8], [8.0], and [24] notation in the remainder of the paper.) was cross-matched with optical photometry ($g,r,i,z,y$) obtained from the Panoramic Survey Telescope and Rapid Response System (Pan-STARRS1; Chambers et al., 2016) or PS1 data archive, which is not available for the most southern galaxies. Additional photometry was obtained from the VISTA Hemisphere Survey (VHS; McMahon 2012), and the UK Infra-Red Telescope (UKIRT) Hemisphere Survey (UHS; Dye et al. 2018; Irwin 2013), with a limited, however, coverage of our sample for each survey [^5] For only one case, NGC 5253, due to the lack of PS1 data and the availability of VHS we used the y-VHS instead..
@@ 2.3 Removing foreground stars
At the start of the project, Gaia DR2 was available. This release included a meaningful number of measurements (for parallax and proper motion) only for a few galaxies. Therefore, we initially performed the foreground selection based on Gaia data for the galaxies M31 and M33, to derive specific cuts on parallax and proper motion (see Paper I). With the release of DR3 (Gaia Collaboration et al., 2023) both the number of available data increased significantly and the quality of the measurements has improved. Therefore, we decided to revisit this process.
@@
We performed foreground removal for each galaxy independently, defining parallax and proper motion limits based solely on sources within each galaxy. For this process we defined an appropriate ellipse around each galaxy (using an optical image of the galaxy from the Digitized Sky Survey (DSS)-red) that would enclose the majority of its sources. This border does not coincide (necessarily) with any other radial profile derived from other means (e.g. $D_{25}$ or Petrosian radius), but it does not affect our results, which are based on statistical treatment of the data. We next defined a search radius for Gaia sources around each galaxy by taking a radius about twice larger than the semi-major axis of the ellipse defined around the galaxy (in contrast to the box defined around M31 and M33 in Paper I - however, the change does not affect the end result). We cross-matched (using a 1″  search radius) the Gaia queried catalog around each galaxy with the corresponding catalog we had built from IR and optical photometry. Most of the galaxies have Gaia data, but this depends on the distance and the crowding for each galaxy (number of Gaia sources are presented in Table 2).
@@
Our purpose is to optimally select the foreground sources within the galaxy so that we eliminate (minimize, in reality) contamination, as foreground and galactic sources cannot be distinguished from the original IR catalogs. To achieve this we followed the Gaia source selection from Lindegren et al. (2021). For this, we first calculate the median error of each of the quantities parallax_error, pmra_error, and pmdec_error, and we used this to replace all corresponding values that were smaller than their median ones (as these might be underestimates of the real error). We excluded all sources without Gaia data, and then we selected sources with the following criteria: astrometric_excess_noise¡1, pmra_error / pmdec_error ¡3, parallax_error¡1.5, and phot_g_mean_mag $\leq 20.7$ (updated for DR3). Following Paper I, we first derived the distributions of the quantities parallax/parallax_error [^6] In Paper I we applied a 0.03 mas offset (Lindegren et al., 2018). However, it seems that its value is smaller and possibly consistent with 0 (Groenewegen, 2021). In any case, the inclusion of this offset would lead to a subtle difference in the parallax/parallax_error cut, which has no real impact in our work, and hence we did not apply it., pmra/pmra_error, and pmdec/pmdec_error for all sources outside the ellipse defined for each galaxy. These foreground distributions were fit with a spline function. Then, we plotted the equivalent distributions for the sources within the ellipse, which consist of the foreground population and the true galactic one. The foreground contamination is estimated by the (previously defined) spline scaled according to the ratio of the density inside and outside the galaxy. For the galactic population, we assumed a Gaussian distribution. The total distribution is therefore fit by combining this Gaussian along with the foreground contamination. After obtaining the mean value and standard deviation of this Gaussian profile, we determined as foreground sources those sources with values larger than $3\sigma$. For the parallax/parallax_error we considered as foreground sources only those with values larger than $+3\sigma$ since negative values are not valid measurements. The selected values for each galaxy are presented in Table 2. We also kept all sources without Gaia data, as there is no way we can decide whether such a source is foreground or galactic. By doing this, we actually increased the contamination but we were certain that we did not exclude possible interesting sources.
@@
This step above was successful for the majority of the galaxies. However, there were cases where either the number of sources within the galaxy was too low (e.g. 17 stars in NGC 3077) or the distribution was very sparse and the fit was not trustworthy (e.g. IC 10 with 376 sources within the galaxy). For all these cases (11 out of 26 galaxies, indicated with an ’*’ in their parallax and proper motion limits in Table 2) we used the criteria derived from M31, which is the most populated galaxy.
@@
During the cross-matching of the original IR catalogs with other surveys (including the Gaia DR2 initially) all multiple matches were removed (typically less than $\sim 3\%$; Paper I). Upon inspection of the results we found that a very small fraction of IR sources matched the same Gaia DR3 source. These sources were present in the initial point-source catalogs with different IDs, but they had similar coordinates and magnitudes. We considered them as duplicates in the initial catalogs. Although they matched the same Gaia source, the different magnitudes in IR could lead to different classification results. To avoid confusion we opted to remove these sources altogether. Depending on the initial IR catalog this fraction was less than $0.1\%$ (e.g. in NGC 4736 and M83 from Khan 2017, or NGC 6822, NGC 300, M81 from Khan et al. 2015), to $\sim 0.3\%$ (e.g. in NGC 253 and NGC 55; Williams & Bonanos 2016) and up to $\sim 7\%$ (e.g. in IC 10, IC 1613, and WLM; Boyer et al. 2015). The final number of sources per galaxy that was analyzed in this work is presented in the last column of Table 2.

### v2

@@ Abstract
Methods. We combined Spitzer photometry and Pan-STARRS1 optical data to classify point sources in 26 galaxies within 5 Mpc, and a metallicity range 0.07-1.36 Z. Gaia data release 3 (DR3) astrometry was used to remove foreground sources. Classifications are derived using a machine-learning model developed in our previous work.
Results. We report classifications for 1,147,650 sources, with 276,657 sources ($\sim 24\%$) being robust. Among these are 120,479 red supergiants (RSGs; $\sim 11\%$). The classifier performs well even at low metallicities ($\sim 0.1$ Z) and distances under 1.5 Mpc, with a slight decrease in accuracy beyond $\sim 3$ Mpc due to Spitzer ’s resolution limits. We also identified 21 luminous RSGs ($\textrm{log}(L/L_{\odot})\,$$\geq 5.5$), 159 dusty yellow hypergiants in M31 and M33, as well as 6 extreme RSGs ($\textrm{log}(L/L_{\odot})\,$$\geq 6$) in M31, challenging observed luminosity limits. Class trends with metallicity align with expectations, although biases exist.
Conclusions. This catalog serves as a valuable resource for individual-object studies and James Webb Space Telescope target selection. It enables the follow-up on luminous RSGs and yellow hypergiants to refine our understanding of their evolutionary pathways. Additionally, we provide the largest spectroscopically confirmed catalog of extragalactic massive stars and candidates to date, beyond the Clouds, comprising 5,273 sources (including $\sim 330$ other objects).
@@
Stars: massive – Stars: mass-loss – Stars: evolution – Methods: statistical – Catalogs
@@ 1 Introduction
One of the main goals of the James Webb Space Telescope (JWST) is to study galaxies over cosmic time, from the early Universe to the present. However, their light is a combination of various components, one of which is their stellar content. Although rare in absolute numbers, massive stars make an important contribution. Through their feedback, whether from strong stellar winds or explosive supernovae, massive stars play a critical role in enriching and shaping the environments of their host galaxies. This is especially important in the early Universe, when metallicity was extremely low; the formation and evolution of such stars is still not well understood. The only way to gain insight into these objects is by examining these populations in nearby low metallicity galaxies. Resolved population studies in such galaxies are possible in the Local Group (e.g., Sextans A; Lorenzo et al. 2022, Large Magellanic Cloud (LMC); Vink et al. 2023, Small Magellanic Cloud (SMC); Shenar et al. 2024) but challenging at longer distances. Therefore, we lack well-explored populations of massive stars at these metallicities.
@@
The main goal of the ASSESS [^1] http://assess.astro.noa.gr/ (Episodic Mass Loss in Evolved Massive stars: Key to Understanding the Explosive Early Universe) project (Bonanos et al., 2024) was to investigate the role of episodic mass loss in the evolution of massive stars (see e.g., Yang et al. 2023; Antoniadis et al. 2024, 2025; de Wit et al. 2024; Munoz-Sanchez et al. 2024a, b; Zapartas et al. 2025; Christodoulou et al. 2025). A large number of sources with secure classifications was needed to explore its importance across various metallicity environments; therefore, we set up both an observing campaign to acquire spectra for a large number of sources (see e.g., de Wit et al., 2023; Bonanos et al., 2024; de Wit et al., 2025) as well as a machine-learning approach (Maravelias et al. 2022; henceforth Paper I). We developed a classifier that uses optical and IR photometry to select dusty, mass-losing, evolved massive stars and classify Spitzer -detected point sources into different broad classes. The purpose of this classifier was to predict the classes for approximately 1.2 M sources from 26 galaxies within 5 Mpc and spanning a metallicity range (0.07-1.36 Z), creating the largest point source catalog with spectral-type classifications for and beyond the Local Group. In Paper I, we provide the method and explore the classifier’s prediction accuracy [^2] Code and other material available at https://github.com/gmaravel/pc4mas., while the current paper presents the results of the classifier’s application and the corresponding catalog of sources.
@@
In Sect. 2 we present the data collection and processing, the selection process of the best candidates (based on the results of the machine-learning classifier), as well as the collection of all sources with known spectral classification from the literature. In Sect. 3, we describe the catalog we compiled, along with statistics on the number of objects per class, and present exemplar color-magnitude diagrams (CMDs). In Sect. 4, we provide a comparison of the spectral types we predicted with those derived from the literature and discuss the performance of the classifier. We present the populations as functions of metallicity and explore the luminous red supergiants (RSGs) and dusty yellow supergiants (YSGs) found in our sample. Finally, in Sect. 5 we summarize and conclude our work.
@@ 2 Data collection and processing
In the following sections, we describe our sample and the steps we followed to build our catalogs and remove foreground sources. Although this has been extensively discussed in Paper I, we provide here a short overview, along with adjustments mainly in the foreground cleaning approach prompted by the release of Gaia data release 3 (DR3).
@@ 2.1 Galaxy sample
In Paper I, we used M31 and M33 to train the classifier, and WLM, Sextans A, and IC 1613 to test it. In this work, we present the results of applying the classifier to the whole sample of 26 galaxies included in the ASSESS project (Bonanos et al., 2024) (except for the Clouds, which were treated separately; e.g., Yang et al. 2019, 2020, 2021, 2023; de Wit et al. 2023; Antoniadis et al. 2024; Munoz-Sanchez et al. 2024a). This is presented in Table 1, along with the basic properties of the galaxies, including their names, coordinates, types, radii (corresponding to galaxy sizes, based on visual inspection, used to match the catalogs), distances, and metallicity.
Table 1: Galaxies examined in this work along with some basic properties.333\$a\$\$a\$footnotetext: Derived from NASA/IPAC Extragalactic Database (NED). \$b\$\$b\$footnotetext: From Tully et al. (2023), except for M81, Sextans B, NGC 3109, and NGC 5253, which were derived from Tully et al. (2013).\$c\$\$c\$footnotetext: Metallicities are based on young clusters or massive stars or H $i$ regions, when available. An average value was reported in cases where multiple measurements or radial gradients exist. For solar abundance we used 12+log(O/H) = 8.69 (Asplund et al., 2009).
@@
($h$:$m$:$s$) ($d$:$m$:$s$) (′) (Mpc) ($Z_{\sun}$)
@@
[…]
1Urbaneja et al. (2008); 2Hartoog et al. (2012); 3Cosens et al. (2024); 4Zurita and Bresolin (2012, range 1.05-1.66 $Z_{\sun}$); 5Davidge 2021, similar to LMC, 0.43 Z from Hunter et al. (2007) and M33 - see below; 6Spinoglio et al. (2022); 7Kudritzki et al. (2008); 8Bresolin et al. (2007); 9U et al. (2009, range 0.3-1.0 $Z_{\sun}$); 10Ross et al. (2015); 11Hadfield and Crowther (2007); 12Thuan and Izotov (2005); 13Bresolin et al. (2022, range 0.35-0.77 $Z_{\sun}$); 14Arellano-Córdova et al. (2016, range 0.37-0.78 $Z_{\sun}$); 15Kniazev et al. (2005, range 0.07-0.14 $Z_{\sun}$); 16Hosek et al. (2014); 17Storchi-Bergmann et al. (1994); 18Pilyugin et al. (2015); 19Moustakas and Kennicutt (2006); 20Kang et al. (2020, from iron abundance, range 0.03-0.40 $Z_{\sun}$); 21Hernandez et al. (2019, range 0.35-1.58 $Z_{\sun}$); 22Monreal-Ibero et al. (2012); 23Patrick et al. (2015); 24Skillman et al. (1997, range 0.06-0.20 $Z_{\sun}$) 25Della Bruna et al. (2021).
Figure 1: Example of fitting processing and foreground removal in NGC 2403. Top panel: Gaia sources where green ellipse defines boundary selected for galaxy. Middle panel: Fitting parallax data for all sources within NGC 2403. The foreground contribution (dotted dark gray line) is scaled according to the densities of sources inside and outside the galaxy, while the dashed purple line shows the (Gaussian) contribution of galactic sources, and the blue line indicates the total fit. The vertical dashed black lines show the $3\sigma$ limits defined by the Gaussian properties of the galactic sources. Bottom panel: Similar to parallax, but for proper motion in declination (see Sect. 2.3 for more details).
@@ 2.2 Surveys used and data processing
Since IR photometry has been proven successful in identifying evolved massive stars (Britavskiy et al., 2014, 2015, 2019; Kourniotis et al., 2017) and those with dusty environments (Bonanos et al., 2009, 2010, 2024), we based our catalogs on precompiled point-source catalogs from the Spitzer Space Telescope (Boyer et al. 2015; Khan et al. 2015; Khan 2017; Williams and Bonanos 2016); (see Table 4 for details; the Spitzer column provides the number of sources within the radius as presented in Table 1). The mid-IR photometry ($\rm 3.6\,\mu m$, $4.5\,\mu\rm m$, $5.8\,\mu\rm m$, $8.0\,\mu\rm m$, $24\,\mu\rm m$ [^4] We are using the [3.6], [4.5], [5.8], [8.0], and [24] notation in the remainder of the paper.) was crossmatched with optical photometry ($g,r,i,z,y$) obtained from the Panoramic Survey Telescope and Rapid Response System (Pan-STARRS1; Chambers et al., 2016), or PS1, data archive, which is not available for the most southern galaxies. Additional photometry was obtained from the VISTA Hemisphere Survey (VHS; McMahon 2012) and the UK Infra-Red Telescope (UKIRT) Hemisphere Survey (UHS; Dye et al. 2018; Irwin 2013), however, with limited coverage of our sample for each survey [^5] For only one case, NGC 5253, due to the lack of PS1 data and the availability of VHS, we used the y-VHS instead..
@@ 2.3 Removing foreground stars
At the start of the project, Gaia data release 2 (DR2) was available. This release included a meaningful number of measurements (for parallax and proper motion) only for a few galaxies. Therefore, we initially performed foreground selection based on Gaia data for galaxies M31 and M33, to derive specific cuts in parallax and proper motion (see Paper I). With the release of DR3 (Gaia Collaboration et al., 2023), both the amount of available data increased significantly and the quality of the measurements improved. Therefore, we decided to revisit this process.
@@
We performed foreground removal for each galaxy independently, defining parallax and proper motion limits based solely on sources within each galaxy. For this process, we defined an appropriate ellipse around each galaxy (by visual inspection of its optical image from the Digitized Sky Survey (DSS; red)) that enclosed the majority of its sources. This approach traces the spatial extent of resolved stellar populations providing a concrete and physically motivated boundary (in contrast to Petrosian and $D_{25}$ radii that are derived from integrated light profiles and brightness thresholds, respectively, and may not capture the true, sometimes fainter, outer edge [^6] However, this approach is limited to galaxies with Gaia data and does not work for galaxies at larger distances for which the Petrosian and $D_{25}$ radii can provide homogeneous results.). Additionally, we are interested in the statistical treatment of the data, so even though we may include larger radii in some cases, we are confident that we always properly sample the foreground population (which, in turn, is scaled for the size of the galaxy; see below). We next defined a search radius for Gaia sources around each galaxy by taking a radius about twice as large as the semi major axis of the ellipse defined around the galaxy (in contrast to the box defined around M31 and M33 in Paper I; however, the change does not affect the end result). We crossmatched (using a 1″  search radius) the Gaia queried catalog around each galaxy with the corresponding catalog we had built from IR and optical photometry. Most galaxies have Gaia data, but this depends on the distance and crowding for each galaxy (the number of Gaia sources is presented in Table 4).
@@
Our purpose is to optimally select the foreground sources within the galaxy so that we eliminate (minimize, in reality) contamination, as foreground and galactic sources cannot be distinguished from the original IR catalogs. To achieve this, we followed the Gaia source selection from Lindegren et al. (2021). For this, we first calculated the median error of each of the quantities parallax_error, pmra_error, and pmdec_error and used this to replace all corresponding values that were smaller than their median values (as these might be underestimates of the real error). We excluded all sources without Gaia data and then selected sources with the following criteria: astrometric_excess_noise¡1, pmra_error / pmdec_error ¡3, parallax_error¡1.5, and phot_g_mean_mag $\leq 20.7$ (updated for DR3). Following Paper I, we first derived the distributions of the quantities parallax/parallax_error [^7] In Paper I we applied a 0.03 mas offset (Lindegren et al., 2018). However, it seems that its value is lower and possibly consistent with 0 (Groenewegen, 2021). In any case, the inclusion of this offset led to a subtle difference in the parallax/parallax_error cut, which has no real impact in our work, and hence we did not apply it., pmra/pmra_error, and pmdec/pmdec_error for all sources outside the ellipse defined for each galaxy. The distribution of the properties of these foreground sources was modeled using a regularized cubic spline (scipy.interpolate.UnivariateSpline, $k=3$) fitted to the histogram bin centers, with the smoothing parameter $s$ empirically adjusted to ensure smoothness without overfitting noise. Then, we plotted the equivalent distributions for the sources within the ellipse, which consist of the foreground population and the true galactic one. The foreground contamination is estimated by the (previously defined) spline and scaled according to the ratio of the density inside and outside the galaxy, assuming that the foreground properties are approximately uniform across the field, as we estimated these properties from a region with a radius at least twice the size of each galaxy. For the galactic population, we assumed a Gaussian distribution. The total distribution was therefore fitted by combining this Gaussian along with the foreground contamination. After obtaining the mean value and standard deviation of this Gaussian profile, we determined as foreground sources those sources with values higher than $3\sigma$. For the parallax/parallax_error we considered as foreground sources only those with values higher than $3\sigma$ since negative values are not valid measurements. The selected values for each galaxy are presented in Table 4. We also kept all sources without Gaia data, as there is no way we can decide whether such a source is foreground or galactic. By doing this, we actually increased contamination, but we were certain that we did not exclude possible interesting sources.
@@
This step above was successful for the majority of the galaxies. However, there were cases where either the number of sources within the galaxy was too low (e.g., 17 stars in NGC 3077) or the distribution was very sparse and the fit was not trustworthy (e.g., IC 10 with 376 sources within the galaxy). For all these cases (11 out of 26 galaxies, indicated with an asterisk in their parallax and proper motion limits in Table 4) we used the criteria derived from M31, which is the most populated galaxy.
@@
During crossmatching of the original IR catalogs with other surveys (including the Gaia DR2 initially), all multiple matches were removed (typically less than $\sim 3\%$; Paper I). Upon inspection of the results, we found that a very small fraction of IR sources matched the same Gaia DR3 source. These sources were present in the initial point source catalogs with different IDs, but they had similar coordinates and magnitudes. We considered them as duplicates in the initial catalogs. Although they matched the same Gaia source, the different magnitudes in IR could lead to different classification results. To avoid confusion, we opted to remove these sources altogether. Depending on the initial IR catalog, this fraction was less than $0.1\%$ (e.g., in NGC 4736 and M83 from Khan 2017, or NGC 6822, NGC 300, M81 from Khan et al. 2015), to $\sim 0.3\%$ (e.g., in NGC 253 and NGC 55; Williams and Bonanos 2016) and up to $\sim 7\%$ (e.g., in IC 10, IC 1613, and WLM; Boyer et al. 2015).